# Evaluating Large Language Model based Personal Information Extraction and Countermeasures

### v1

@@ 9 Conclusion, Limitations, and Future Work
Limitation and future work: In this work, we propose to utilize prompt injection to protect personal information against LLM-based PIE. We extend several existing techniques against prompt injection as adaptive attacks. Our results show they achieve limited effectiveness. However, we note that an attacker may develop more advanced techniques against prompt injection to bypass prompt injection-based protection. As our work aims to perform a measurement study on existing attacks and defenses, we leave the development of more advanced techniques against prompt injection as an interesting future work.

### v2

@@ 10 Conclusion, Limitations, and Future Work
Limitation and future work: In this work, we propose to utilize prompt injection to protect personal information against LLM-based PIE. We extend several existing techniques against prompt injection as adaptive attacks. Our results show they achieve limited effectiveness. However, we note that an attacker may develop more advanced techniques against prompt injection to bypass prompt injection-based protection. As our work aims to perform a measurement study on existing attacks and defenses, we leave developing more advanced techniques against prompt injection as future work.
11 Ethics Considerations
Our research centers on evaluating LLM-based personal information extraction and corresponding countermeasures. The primary aim is to enhance the community’s understanding of LLM-based personal information extraction and to encourage the development of more robust mitigation strategies. By identifying potential vulnerabilities and offering solutions, this research makes a positive contribution to AI security, promoting the protection of personal information.
12 Open Science
Our code and Synthetic dataset are available at https://doi.org/10.5281/zenodo.14737200. We acknowledge the dual-use nature of releasing code and datasets related to this research, which has the potential to assist both defenders and attackers. To address potential ethical concerns about open science compliance, we will share the other datasets only with verified researchers upon request, rather than releasing them publicly. Researchers requesting access will be required to provide institutional email addresses, professional affiliations, and a brief explanation of how the code and data will be used. This ensures that the code and datasets are shared responsibly and minimizes the risk of misuse while still enabling replication and further research.
Acknowledgment:  We thank reviewers and shepherd for constructive feedback. This work was supported by NSF under grant No. 2414407, 2131859, 2125977, 2112562, 1937787, ARO under grant No. W911NF2110182, as well as Microsoft Accelerating Foundation Models Research program.